OpenStudy (anonymous):

Which of the following values in the set below will make the equation 6x + 2 = 14 true? (Only input the number.) {0, 1, 2, 3, 4}

OpenStudy (anonymous):

Help

OpenStudy (anonymous):

Answer Please

OpenStudy (anonymous):

Hello????

OpenStudy (texaschic101):

you could just do some subbing of your answers into the equation to see what is true....or you can just solve the equation. 6x + 2 = 14 --- subtract 2 from both sides 6x = 14 - 2 6x = 12 -- now divide both sides by 6 x = 2 understand ?? Any questions ?

OpenStudy (anonymous):

I still don't get it

OpenStudy (texaschic101):

6x + 2 = 14.....we want to get the x term on one side of the equal sign, and everything else on the other side. So we will subtract 2 from both sides. 6x + 2 - 2 = 14 - 2 (and what you do on one side of the equal sign, you have to do on the other, Now we will combine terms.. 6x = 12....you see how the +2 and the -2 on the left side cancel each other out. Now we must divide both sides by 6 to get x by itself (6/6)x = 12/6 --- 6/6 reduces to 1 and 1x is the same as x....so basically, we are left with x = 2 does that help ?

OpenStudy (anonymous):

Thank you

OpenStudy (texaschic101):

your welcome :)